the manager of an italian restaurant is considering a carryout food service. however the manager is concerned that not all individuals placing orders by ohone actually pick up order. if 90% or less of phone orders are picked up, it will not have a profitable carry out operation. it can be concluded that more than 90% will have a profitable income. during two week test period, 234 orders in a sample of 250 phone in orders were picked up. using a 0.05 level of significance what conclusion should be made?
so i went ahead and calculated the Z(alpha/2) value and i got 1.96. i also calculated the proportion which had a value of 0.936 (which is represented by p) then i calculated q which was 0.064. i dont know what to do now because the question already gives the interval so what exactly am i supposed to conclude if the interval is already explaining that over 90% is profitable and less than 90% is nnot.

Answer by ewatrrr(24785)

Ho: p ≤ 0.90
Ha: p > .90 (claim)
.........
Sample: p = 234/250 = .936
.....
z (.936) = (.936-.90)/sqrt(.90*.10/250)) = 1.897
P( z > 1.897) = .0289
......
95% confidence level
.0289 < .05
.........
Ho rejected, claim p > .90 accepted